OpenStudy (anonymous):

I need help to solve this equation please; g(x) = log(x +5)

OpenStudy (anonymous):

there is nothing to solve here. this is just the definition of a function

OpenStudy (anonymous):

Thank you but what does that mean

OpenStudy (anonymous):

it means take your input x, add 5 then take the common log (log base ten)

OpenStudy (anonymous):

so for example \[g(95)=\log(95+5)=\log(100)=2\]

OpenStudy (anonymous):

You probably are being asked to graph this function. I will attach the graph. The graph intercepts the x-axis at the point (-4,0); it intercepts the y-axis at (0, log 5). There is a vertical asymptote with equation x=-5 (indicated by the dashed line on the graph). The domain is x>-5, and the range is all y real numbers.